OpenStudy (anonymous):

0.16x-0.53=1.48x-.77x-3.527

OpenStudy (anonymous):

3.527-0.53 = 1.48 - 0.77x - 0.16x 2.997 = 0.55x 5.449.. = x

OpenStudy (anonymous):

so the answer is 5.449?

OpenStudy (anonymous):

actually its a fraction.. so not exactly do it in your calculator to find out

OpenStudy (anonymous):

ok